OpenStudy (anonymous):

write the expression 64^-1/3

OpenStudy (jdoe0001):

keep in mind that \(\bf \large a^{-\frac{{\color{blue} n}}{{\color{red} m}}} = \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}} \implies \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}}\)

OpenStudy (anonymous):

write an expression 64^-1/3 in radical notation

OpenStudy (isaiah.feynman):

Use jdoe explanation to guide yourself.

OpenStudy (isaiah.feynman):